# Meritocracy and Its Discontents: Long-run Effects of Repeated School Admission Reforms

### v1

@@ A. Baseline Specification
Tokyo area ($<$ 100 km) 0.19 -0.02 0.07 0.17 0.11 -0.01 0.02
  $\times$ Time trend (0.673) (0.802) (0.822) (0.476) (0.760) (0.879) (0.835)
@@ B. Adding Control Variables
Tokyo area ($<$ 100 km) 0.21 -0.01 0.10 0.18 0.14 -0.01 0.02
  $\times$ Time trend (0.630) (0.836) (0.751) (0.448) (0.701) (0.885) (0.841)
@@ C. Squared trend term
Tokyo area ($<$ 100km) $\times$ Time trend -0.93 -0.08 -0.48 0.02 0.09 -0.13 0.22
(0.415) (0.817) (0.275) (0.981) (0.903) (0.629) (0.412)
[…]
(0.465) (0.824) (0.288) (0.878) (0.927) (0.695) (0.333)
Observations 470 470 470 470 470 470 470
Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
Mean dep var 5.02 1.21 4.50 4.34 6.34 0.66 0.39
Mean dep var (Tokyo area 6.47 1.29 5.96 5.23 8.80 0.61 0.54
@@
(1) (2) (3) (4) (5) (6) (7)
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
Univ. income income recipients executives politicians & Univ.
grads earners earners bureaucrats professors
Age 17 under centralization 1.71 0.50 1.81 1.44 1.11 0.39 0.08
  $\times$ Tokyo area ($<$100 km) (0.016)** (0.044)** (0.029)** (0.070)* (0.013)** (0.018)** (0.717)
[0.004]*** [0.003]*** [0.004]*** [0.008]*** [0.008]*** [0.282] [0.496]
Observations 705 705 705 705 705 705 705
Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
Mean dep var 4.79 0.81 3.64 3.86 2.80 1.18 0.70
Mean dep var (Tokyo area 5.68 1.11 6.18 4.46 4.46 1.14 0.70
@@ B. Adding Control Variables
Mean dep var 11.67 0.94 0.29 297.92
Mean dep var (Tokyo area 13.18 2.89 0.37 463.52
@@
Notes: This figure plots the average number of years taken from entering a higher school to graduating from an Imperial University by cohort (where cohort is defined by the year of university graduation) for two groups of individuals who passed the administrative division of the Higher Civil Service Exams. The first group is top-ranking officials who graduated from Schools 1–8 (shown in black round markers). The second group is non-top-ranking officials who graduated from Schools 1–8 (shown in red rectangular markers for two representative years, 1914 and 1922, only). For both groups, we find the exact year of entering Schools 1–8 for each individual using the Student Registers. The figure shows that the average number of years is not significantly different between the two groups. See Section 4.2 for discussions about this figure.
@@
Figure A.10: Share of Government Officials in PIR-listed Individuals
Notes: This figure plots the share of central government officials in all individuals listed in PIR (1939) by birth cohort. The cohorts are colored according to their exposure to the centralized admissions in 1902–1907, where darker color indicates higher intensity of exposure. We observe no significant increase in the share of central government officials for the cohorts who are exposed to the centralized admissions. See Section 4.2 for discussions about this table.
Figure A.11: Total Number of PIR-listed Individuals by Cohort
Notes: This figure plots the total number of individuals listed in PIR (1939) and PIR (1934) by birth cohort. The cohorts are colored according to their exposure to the centralized admissions in 1902–1907, where the darker color indicates the higher intensity of exposure. In both PIR (1939) and PIR (1934), cohorts exposed to the centralization tend to have a greater number of individuals selected into the PIR. See Appendix Table A.17 for a statistical analysis and Section 4.2 for discussions about this figure.
Table A.17: Long-run Impacts: Total Number of PIR-listed Individuals
(1) (2)
No. of Individuals in PIR
(No. of individuals aged 50 in PIR = 100)
Age 17 under centralization 3.77
(0.086)*
Cohort’s exposure to centralization 4.78
(0.060)*
Age 40.32 39.69
(0.000)*** (0.000)***
Age squared -0.39 -0.38
(0.000)*** (0.000)***
PIR (1939) dummy 0.37 0.37
(0.808) (0.805)
Observations 46 46
Notes: This table examines the long-run effects of the centralized admissions on the total number of career elites in the PIR. To distinguish the cohort effect from the age effect, we use two editions of the PIR, i.e., PIR (1934) and PIR (1939), count the number of individuals aged 40–62 by birth cohort in each edition (see Appendix Figure A.11), and pool these data. We standardize the dependent variable by setting the number of individuals at age 50 in each edition to be 100. “Age 17 under centralization” takes 1 if the cohort turned 17 years old under the centralized system in 1902–1907, and takes 0 otherwise. “Cohort’s exposure to centralization” is defined as in Table 2 Panel D. We control for quadratic age trends and the edition fixed effect. Parentheses contain p-values based on robust standard errors. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 4.2 for discussions about this table.
@@ Appendix B Online Theoretical Appendix
To guide our empirical investigation, we develop a simple model to predict the impacts of centralization on application behavior and assignment. We first confirm that centralized admissions (Capp) was indeed designed to make the school seat allocation more meritocratic compared to decentralized admissions (Dapp). Our model also predicts changes in application behavior. First, a greater number of applicants apply to the most popular school under Capp than under Dapp (Proposition 1). Second, when high-achieving applicants are disproportionately located in a geographical area (the urban area for example), then a larger fraction of school seats are assigned to students coming from that advantageous area (Proposition 2). Finally, a smaller fraction of applicants are assigned to their local school under Capp relative to Dapp (Proposition 3).

### v2

@@ A. Baseline Specification
Tokyo area ($<$ 100 km) 0.19 -0.02 0.07 0.17 0.11 -0.01 0.02 0.92
     $\times$ Time trend (0.673) (0.802) (0.822) (0.476) (0.768) (0.879) (0.835) (0.511)
@@ B. Adding Control Variables
Tokyo area ($<$ 100 km) 0.21 -0.01 0.10 0.18 0.13 -0.01 0.02 1.01
    $\times$ Time trend (0.630) (0.836) (0.751) (0.448) (0.703) (0.885) (0.841) (0.463)
@@ C. Squared Trend Term
Tokyo area ($<$ 100km) -0.93 -0.08 -0.48 0.01 0.06 -0.13 0.22 -0.11
    $\times$ Time trend (0.415) (0.817) (0.275) (0.990) (0.927) (0.629) (0.412) (0.923)
Tokyo area ($<$ 100km) 0.07 0.00 0.03 0.01 0.00 0.01 -0.01 0.07
    $\times$ Time trend2 (0.465) (0.824) (0.288) (0.870) (0.893) (0.695) (0.333) (0.586)
Observations 470 470 470 470 470 470 470 470
Cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes Yes
Mean dep var 5.02 1.21 4.50 4.34 6.71 0.66 0.39 28.62
Mean dep var (Tokyo area 6.47 1.29 5.96 5.23 9.21 0.61 0.54 38.97
@@
$Y_{pt}=\beta\times Trend_{t}\times Tokyo\_area_{p}+\alpha_{p}+\alpha_{t}+\epsilon_{pt},$
@@
$Y_{pt}=\beta_{1}\times Trend_{t}\times Tokyo\_area_{p}+\beta_{2}\times(Trend_{t})^{2}\times Tokyo\_area_{p}+\alpha_{p}+\alpha_{t}+\epsilon_{pt},$
@@
(1) (2) (3) (4) (5) (6) (7) (8)
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial All
Univ. income income recipients executives politicians & Univ. occupational
grads earners earners bureaucrats professors elites
Age 17 under centralization 1.71 0.50 1.81 1.44 1.20 0.39 0.08 4.86
(0.016)** (0.044)** (0.029)** (0.070)* (0.026)** (0.018)** (0.717) (0.072)*
[0.011]** [0.004]*** [0.005]*** [0.002]*** [0.001]*** [0.261] [0.690] [0.005]***
Observations 705 705 705 705 705 705 705 705
Cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes Yes
Mean dep var 4.79 0.81 3.64 3.86 2.95 1.15 0.52 14.91
Mean dep var (Tokyo area 5.68 1.11 6.17 4.46 4.63 1.13 0.70 22.35
@@ B. Adding Control Variables
Mean dep var 11.67 0.93 0.29 297.92
Mean dep var (Tokyo area 13.18 2.88 0.37 463.52
@@
Notes: This figure plots the average number of years taken from entering a higher school to graduating from an Imperial University by cohort (where cohort is defined by the year of university graduation) for two groups of individuals who passed the administrative division of the Higher Civil Service Exams. The first group is top-ranking officials who graduated from Schools 1–8 (shown in black circles). The second group is non-top-ranking officials who graduated from Schools 1–8 (shown in red rectangular markers for two representative years, 1914 and 1922, only). For both groups, we find the exact year of entering Schools 1–8 for each individual using the Student Registers. The figure shows that the average number of years is not significantly different between the two groups. See Section 4.2 for discussions about this figure.
@@
Figure A.10: Share of Government Officials in Elites Listed in the PIR
Notes: This figure plots the share of central government officials in all individuals listed in the PIR (1939) by birth cohort. The cohorts are colored according to their exposure to the centralized admissions in 1902–1907, where darker color indicates higher intensity of exposure. We observe no significant increase in the share of central government officials for the cohorts exposed to the centralized admissions. See Section 4.2 for discussions about this table.
Figure A.11: Total Number of All Occupational Elites by Cohort
Notes: This figure plots the total number of all occupational elites listed in the PIR (1939) and PIR (1934) by birth cohort. The cohorts are colored according to their exposure to the centralized admissions in 1902–1907, where the darker color indicates the higher intensity of exposure. In both PIR (1939) and PIR (1934), cohorts exposed to centralization tend to have a greater number of individuals selected into the PIR. See Table 5 for a statistical analysis and Section 4.2 for discussions about this figure.
@@ Appendix B Online Theoretical Appendix
To guide our empirical investigation, we develop a simple model to predict the impacts of centralization on application behavior and assignment. We first confirm that the centralized admissions (referred to as “Capp” in this section) was indeed designed to make the school seat allocation more meritocratic compared to the decentralized admissions (referred to as “Dapp” in this section). Our model also predicts changes in application behavior. First, a greater number of applicants apply to the most popular school under Capp than under Dapp (Proposition 1). Second, when high-achieving applicants are disproportionately located in a geographical area (the urban area for example), then a larger fraction of school seats are assigned to students coming from that advantageous area (Proposition 2). Finally, a smaller fraction of applicants are assigned to their local school under Capp relative to Dapp (Proposition 3).